In an economy labour supply: w=20+4L, labour demand: w=50-2L. Labour is organized in a Union acting as a monopolist. Find the level of labour employed in the economy and the unemployed one
1
Expert's answer
2014-05-27T08:32:03-0400
Labour supply: w=20+4L, labour demand: w=50-2L.Labour is organized in a Union acting as a monopolist.
We should find the level of labour employed in the economy in the point where labour supply equals labour demand, so:
20 + 4L = 50 - 2L
6L = 30,
L = 5 workers.

Their wage is w = 20 + 4*5 = $40
To find the number of unemployed workers we should know the total number of population.